if the hypotenuse is 150cm and the ration for the sides is 16:9. whats are the side lengths and ow do you figure it out?

Answer by rothauserc(4718)

we start with Pythagorean theorem, a^2 + b^2 = h^2
we are given the hypotenuse, 150, and the ratio of the other two sides, 16:9
we can now rewrite the Pythagorean theorem formula for this triangle,
150^2 = (16x)^2 + (9x)^2
22500 = 256x^2 + 81x^2
22500 = 337x^2
x = 8.17 and
side a = 16 * 8.17 = 130.72 cm
side b = 9 * 8.17 = 73.53 cm